OpenStudy (anonymous):

Determine the x-intercept and y-intercept of the following equation. Show all work. 4x - 3y = -24

OpenStudy (texaschic101):

to find the x intercept, sub in 0 for y 4x - 3(0) = -24 4x = -24 x = -24/4 x = - 6 to find the y intercept, sub in 0 for x 4(0) - 3y = -24 -3y = -24 y = -24/-3 y = 8 x intercept is (-6,0) and y intercept is (0,8)

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

no problem :)